# Articulated Object Manipulation using Online Axis Estimation with SAM2-Based Tracking

### v1

@@ Abstract
Articulated object manipulation requires precise object interaction, where the object’s axis must be carefully considered. Previous research employed interactive perception for manipulating articulated objects, but typically, open-loop approaches often suffer from overlooking the interaction dynamics. To address this limitation, we present a closed-loop pipeline integrating interactive perception with online axis estimation from segmented 3D point clouds. Our method leverages any interactive perception technique as a foundation for interactive perception, inducing slight object movement to generate point cloud frames of the evolving dynamic scene. These point clouds are then segmented using Segment Anything Model 2 (SAM2), after which the moving part of the object is masked for accurate motion online axis estimation, guiding subsequent robotic actions. Our approach significantly enhances the precision and efficiency of manipulation tasks involving articulated objects. Experiments in simulated environments demonstrate that our method outperforms baseline approaches, especially in tasks that demand precise axis-based control.
@@ I INTRODUCTION
Robotic manipulation has a wide range of applications, such as industrial automation, medical surgery, and warehouse logistics. Among various manipulation tasks, those involving articulated objects, such as doors and drawers, are particularly challenging. This is because manipulating articulated objects requires not only a deep understanding of their overall geometry but also knowledge of their part compositions and the kinematic relationships between those parts [1].
@@
Traditional manipulation techniques, which typically rely on predefined kinematic models [2, 3] or open-loop control [4], have shown limitations in dynamically adapting to real-world interactions, especially in the presence of uncertainties. One significant drawback of open-loop approaches is their lack of feedback regulation, preventing the robot from adjusting its actions based on real-time interaction dynamics. Without continuous feedback, the system cannot adapt to
@@
Figure 1: Our approach introduces a closed-loop pipeline integrating SAM2 to track the articulated object throughout the manipulation process captured by an RGB-D camera. The mask is subsequently used to segment out the point cloud of the articulated object, followed by explicitly estimating the joint axis using the oriented bounding box (OBB) of the motion part derived from the point cloud. The robot’s action is instructed under the guidance of the motion axis to perform further manipulation for the next loop.
changes in the object’s state, leading to inefficiencies and inaccuracies in manipulation tasks.
@@
Recently, interactive perception has emerged as a promising approach to address these challenges [1, 5, 4, 6]. By actively interacting with the object, robots can gather real-time sensory data that provides insights into the object’s structure and kinematics. Interactive perception aims to bridge the gap between perception and action by using sensory feedback from these interactions. While this method is effective in providing information about the object’s structure and state, it lacks a critical component — the evolving dynamic interactions between the robot and the articulated object over time. This omission limits the robot’s ability to adapt its manipulation policy in real time as the object’s state changes, leading to inefficiencies in tasks that require precise control. To address this limitation, in this paper, we propose a novel closed-loop pipeline that integrates interactive perception with online refined axis estimation, providing the robot with helpful guidance for axis-aware manipulation. The basic idea is illustrated in Fig. 1. Our approach builds on interactive perception and enhances it by using the guidance of an online refined motion axis, which is derived from segmented 3D point clouds. Specifically, we use any interactive perception technique, such as RGBManip [5], leveraging it to induce slight object movements and generate dynamic 3D point cloud frames. These frames are then processed using an advanced segmentation pipeline, Grounding DINO [7] as object detector and the Segment Anything Model 2 (SAM2) [8] for object segmentation, which identifies and isolates the articulated object. By masking the moving components of the object, we can explicitly calculate the motion axis, which in turn informs the robot’s subsequent manipulation policy. In contrast to previous studies, our closed-loop integration of interactive perception with online refined axis estimation allows for more precise and adaptive control, addressing the limitations of traditional methods that operate in an open-loop fashion. By continuously updating the robot’s understanding of the object’s kinematic state, our approach ensures that the robot maintains effective control throughout the manipulation task. We validate our approach through tasks that require precise axis-based control in simulated environments, demonstrating that our method outperforms baseline approaches. Experimental results demonstrate that our pipeline significantly enhances the precision of manipulation tasks involving articulated objects, while strong generalization capability across articulated object manipulation tasks.
@@
    Closed-loop integration of interactive perception and real-time axis estimation: By continuously refining the motion axis based on updates from 3D point clouds, our approach enables adaptive and precise control during manipulation tasks, overcoming the limitations of the open-loop methods.
@@
    Utilization of advanced segmentation models for precise articulation recognition: Our pipeline employs Grounding DINO for object detection and SAM2 for accurate segmentation, allowing the robot to isolate moving components and calculate the motion axis of moving parts of articulated objects in real-time.
@@
    Enhanced generalization and precision in articulated object manipulation: Our method significantly improves manipulation precision and generalization across diverse cabinets in door-opening and drawer-opening tasks, outperforming state-of-the-art baselines and providing consistent axis-aware manipulation.
@@ II-A Vision-based Robotic Manipulation
Visual perception is crucial for robotic manipulation, whose quality heavily depends on the visual modalities. Existing works on vision-based robotic manipulation have adopted different types of visual input, with various visual modalities demonstrating strong potential in different scenarios, while also exhibiting limitations specific to their applications.
RGB images, despite their richness in texture, lack depth information, which is critical for object manipulation and grasping tasks. To address this, UMPNet [9] and Grasp Proposal Networks [10] turned to depth sensors and combined RGB images with depth maps. RGB-D information provides a more direct representation of the 3D world, which, however, can be noisy and affected by transparent and reflective surfaces [11].
To circumvent the above challenge, unlike the recent approaches that switch between different visual modalities [12, 13, 14], RGBManip utilizes RGB-only visual input and directly estimates the 6D pose of objects from multi-view RGB images. It relies on a single RGB camera mounted on its end-effector as its primary visual input. As a result, during the pre-grasping phase, the camera predominantly captures only the immediate vicinity of the handle, inevitably missing crucial information about the interaction dynamics that occur throughout the manipulation process, which is critical for the robot to understand how articulated objects should be manipulated.
On the other hand, point-cloud-based methods have also demonstrated promising performance in vision-based robotic manipulation. Several studies such as Where2Act [15] and SAGCI-System [16] have integrated point cloud information to predict action affordances. Both methods are particularly well-suited for handling articulated objects, but RLAfford [10] emphasizes interaction dynamics which helps to predict manipulation policies effectively. Furthermore, Flowbot3D [17] leverages point cloud motion flows to predict how objects should be manipulated, whose strength is in handling dynamic articulated objects.
Our approach strives to integrate the strengths of RGB-based methods and point cloud-based techniques by tracking RGB inputs and segmenting the point clouds corresponding to the motion part of articulated objects. In order to address the limitation of opened-loop methods, we integrate an additional RGB-D camera into the scene, specifically aimed at capturing the manipulation process.
II-B 3D Point Cloud Segmentation Powered by 2D Foundation Models
Segmentation techniques for 2D images have reached a high level of sophistication, with models such as Segment Anything Model (SAM) [18] standing out as prominent examples. However, the limitations of segmentation methods for 3D point clouds have hindered the development of their application in robotic manipulation.
DeepLabv3 [19] has adapted 2D image segmentation techniques to the 3D point cloud domain by projecting 3D data onto 2D planes and performing segmentation within the transformed space. It exhibits robustness when handling large-scale data, but the projection from 3D point clouds onto 2D planes can lead to a loss of information, which in turn compromises segmentation accuracy.
To overcome this challenge, researchers have proposed some SAM-based methods for 3D point cloud segmentation. Point-SAM [20] extended SAM to 3D domain, which can
@@
generate segmentation masks in response to point and mask prompts. For more precise segmentation such as part-level segmentation, works like PartSLIP [21] and its enhanced version PartSLIP++ [22] perform 3D point cloud segmentation by employing GLIP [23] and SAM on multiple views of a dense point cloud. All these approaches offer different solutions to the obstacle posed by the irregularity scale and complexity of 3D data, while they are also constrained by the limited viewpoints and sparse point clouds in real-world scenarios.
In practice, sparse viewpoint observations are often sufficient to meet the demands of vision-based robotic manipulation, as only a coarse segmentation of the relevant parts of the target objects is typically required. Our approach builds upon the RGBManip environment by integrating a single RGB-D camera and applying noise-robust techniques to process the point cloud, which enhances the feasibility and portability of our method in real-world scenarios.
Previous SAM-based 3D point cloud segmentation techniques have shown promising results. Here, we highlight that for dynamic scenes, leveraging Segment Anything Model 2 (SAM2) as the foundation model for 3D point cloud segmentation offers superior preservation of the scene’s dynamic information. By masking the moving parts of the target objects in the point cloud, our approach captures rich motion-related information that can better inform the manipulation policy.
@@ II-C Axis Estimation for Articulated Objects
However, previous studies for axis estimation are typically open-loop, which causes a loss of the interaction dynamics. To address this issue, we combine interactive perception techniques with axis estimation, introducing a novel closed-loop pipeline. With the moving components of the articulated objects segmented from each frame scene point cloud, we explicitly estimate the axis of articulation by leveraging task-based priors related to the geometric properties of articulated object motion. Our approach does not depend on high-quality point clouds or extensive manipulation, making it promising for application in complex real-world environments.
@@ III METHOD
We introduce a novel approach for manipulating articulated objects, guided by axis estimation derived from SAM2-based tracking. As depicted in Fig. 2, our pipeline consists of 3 core modules: (1) Interactive Perception & Init-Manipulation Module, (2) Tracking & Segmentation Module and (3) Axis Estimation & Manipulation Module.
Initially, the robot applies any interactive perception methods to manipulate the articulated object, creating a slight displacement. Then, using a SAM2-based approach, the moving part of the object’s point cloud in each frame is captured, followed by further manipulation guided by the estimated axis.
@@ III-A Interactive Perception & Init-Manipulation Module
The Interactive Perception & Init-Manipulation Module employs flexible interactive perception methods to initiate manipulation of articulated objects, producing slight displacements for further analysis. Our method incorporates the state-of-the-art RGBManip as the default implementation for this module, chosen for its recent success in interactive perception. However, the module allows for substitution with other interactive perception techniques that may be more suited to specific scenarios, providing adaptability based on application requirements. As RGBManip initiates the manipulation of articulated objects, the camera periodically records dynamic RGB-D data. The captured RGB data will be leveraged for subsequent tracking and segmentation tasks, while the corresponding depth data facilitates the reconstruction of the 3D point cloud of the evolving dynamic scene. This setup ensures precise monitoring of object motion and enhances the fidelity of axis estimation and further manipulation steps.
III-B Tracking & Segmentation Module
The Tracking & Segmentation Module plays a pivotal role in accurately tracking and segmenting the point clouds of dynamic scenes obtained from the Interactive Perception & Init-manipulation Module. Specifically, we exploit Grounding DINO on the initial RGB frame to generate an anchor box for the target object using the textual prompt. The first RGB frame serves as the starting point for SAM2’s tracking process. The anchor box, corresponding to the target object, is input into SAM2 as a box prompt, enabling it to continuously track the object and provide a mask for it throughout each moment of the dynamic manipulation process. With the object’s mask, we can extract the corresponding portion of the 3D point cloud representing the object from the dynamic scene. After a basic noise filtering step, we compute the oriented bounding box (OBB) for the entire cabinet in the initial frame. Subsequently, by subtracting this OBB from each frame’s point cloud and applying another round of noise reduction, we obtain the point cloud representing the moving parts of the articulated objects. For implementation details, the segmentation can be refined by removing the protruding handle from the cabinet’s OBB to obtain a tighter OBB of the cabinet’s body, which can lead to more precise motion-part segmentation.
@@ III-C Axis Estimation & Manipulation Module
The Axis Estimation & Manipulation Module calculates the axis of motion based on the segmented point cloud representing the articulated object’s moving parts identified by the Tracking & Segmentation Module. Task-specific geometric priors play an essential role in achieving accurate axis estimation, i.e., for a prismatic joint, the moving parts translate along the axis, whereas for a revolute joint, the moving parts rotate around the axis. In our simulation settings, the drawer is designed to translate parallel to the xy-plane, while the door rotates about a hinge aligned with the z-axis. By leveraging these task-based priors, we are able to explicitly calculate the joint’s axis. Consider a single action of robotic manipulation, where the OBBs of the initial and final point cloud of the motion components of the articulated object are denoted as $obb_{st}$ and $obb_{ed}$, with their centers designated as $O_{st}$ and $O_{ed}$ respectively. On the one hand, for a prismatic joint, the axis pivot point is defined as the center of $obb_{st}$, and the axis direction is estimated as the direction $\overrightarrow{d}=\overrightarrow{O_{st}O_{ed}}$. On the other hand, for a revolute joint, the axis pivot point is derived by identifying the intersection of the mid perpendiculars along the longer edges in the top-down view of $obb_{st}$ and $obb_{ed}$. With the axis point $P$ established, the axis direction is ascertained by evaluating the sign of the dot product $\overrightarrow{d}\cdot\overrightarrow{t}$, where $\overrightarrow{t}=\overrightarrow{O_{st}P}\times\overrightarrow{z}$, among which $\overrightarrow{z}$ represents the positive direction of the z-axis.
We propose a closed-loop axis estimation refinement policy predicated on a straightforward observation: as the manipulation of the joint assembly increases, the point cloud of its moving component becomes progressively more amenable to accurate reconstruction. The enhancement in the fidelity of the point cloud data, in turn, furnishes the axis estimation process with inputs that are increasingly precise and dependable. For implementation, after each manipulation, we invoke the Tracking & Segmentation Module and the Axis Estimation Module to ascertain the current axis estimation. The Manipulation Module then guides the robot’s subsequent actions according to this estimation. The interactive process reiterates until the robot has executed all designated actions, with the axis estimation being continuously refined throughout the procedure.

### v2

@@ Abstract
Articulated object manipulation requires precise modeling, where understanding the 3D motion constraints of individual articulated components is crucial. Prior research has leveraged interactive perception to facilitate manipulation; however, open-loop approaches often fail to account for the interaction dynamics, limiting their effectiveness. To overcome this limitation, we propose a closed-loop pipeline that integrates interactive perception with online axis estimation and optimization. Our method processes sequentially segmented 3D point clouds of manipulated object parts, continuously refining the motion model throughout the manipulation process. Specifically, our approach builds upon any interactive perception technique, inducing slight object movements to capture point cloud frames of the evolving dynamic scene. These point clouds are segmented using the Segment Anything Model 2 (SAM2) [Ravi et al.], where the moving object part is accurately masked for online axis estimation. The estimated axis then guides subsequent robotic motion planning, ensuring precise and adaptive manipulation. Experimental results in simulated environments demonstrate that our method outperforms baseline approaches, particularly in tasks requiring precise axis-based control. By integrating real-time perception with online optimization, our framework significantly improves both the accuracy and efficiency of articulated object manipulation.
@@ I INTRODUCTION
Robotic manipulation has a wide range of applications, including industrial automation [1], medical surgery [2], and warehouse logistics [3]. Among various manipulation tasks, those involving articulated objects, such as doors and drawers, pose significant challenges [4]. This complexity arises from the need to understand not only the overall geometry of the object but also the composition and kinematic relationships between its articulated components [5].
@@
Traditional manipulation techniques often rely on predefined kinematic models [6, 7] or open-loop control [8, 9]. However, these approaches struggle to dynamically adapt to real-world interactions, particularly in the presence of uncertainties. A key limitation of open-loop methods is the absence of feedback regulation, preventing the robot from
@@
Figure 1: We introduces a closed-loop pipeline integrating SAM2 to track the articulated object throughout the manipulation process captured by an RGB-D camera. The masks are subsequently used to segment out the point cloud of the articulated object, followed by explicitly estimating the joint axis using the oriented bounding boxes (OBBs) of the motion part derived from the point cloud. The robot’s action is instructed under the guidance of the motion axis to perform further manipulation for the next loop.
adjusting its actions based on real-time interaction dynamics. Without continuous feedback, the system fails to accommodate changes in the object’s state, resulting in inefficiencies and inaccuracies in manipulation tasks [10].
Recently, interactive perception has emerged as a promising approach to address these challenges [5, 9, 11, 12]. Interactive perception aims to bridge the gap between perception and action by using sensory feedback from these interactions. By actively interacting with the object, robots can gather real-time sensory data that provides insights into the structure and kinematics of objects. While these methods effectively provide information about the object’s state, it lacks a critical component — the evolving dynamic interactions between the robot and the articulated object over time. This omission limits the robot’s ability to adapt its manipulation policy in real time as the object’s state changes, leading to inefficiencies in tasks that require precise control.
@@
To address this limitation, in this paper, we propose a novel closed-loop pipeline that integrates interactive perception with online refined axis estimation, providing the robot with helpful guidance for axis-aware manipulation. The basic idea is illustrated in Fig. 1. Our approach builds on interactive perception, and enhances it under the guidance of an online refined motion axis, which is derived from segmented 3D point clouds. Specifically, we leverage any interactive perception technique, such as RGBManip [12], as a foundation method to induce slight object movements and generate dynamic 3D point cloud frames. These frames are then processed using an advanced segmentation pipeline, with Grounding DINO [13] as object detector and Segment Anything Model 2 (SAM2) [14] for object segmentation, which identifies and isolates the point cloud of articulated object. By masking out the moving components of the object, we can explicitly calculate the motion axis with the oriented bounding boxes (OBBs), which in turn informs the robot’s subsequent manipulation action. In contrast to previous studies, our closed-loop integration of interactive perception with online refined axis estimation allows for more precise and adaptive control, addressing the limitations of traditional methods that operate in an open-loop fashion. By continuously updating the robot’s understanding of the object’s kinematic state, our approach ensures that the robot maintains effective control throughout the manipulation task. We validate our approach through tasks that require precise axis-based control, such as drawer-opening and door-opening. Simulated results demonstrate that our method significantly outperforms pure interactive perception techniques, enhancing the precision of manipulation tasks involving articulated objects, and shows strong generalization across different articulated objects.
@@
    Utilization of 2D foundation models for 3D-based manipulation. Our pipeline integrates advanced 2D vision techniques, including Grounding DINO for object detection and SAM2 for video segmentation and tracking, and extends their capabilities to 3D point cloud segmentation and tracking. This enables object-level iterative dynamics modeling of articulated objects, allowing for precise and continuous tracking of their motion throughout the manipulation process.
@@
    An online motion axis estimation method based on part-level iterative dynamics. We derive the moving components of the articulated objects from the dynamic evolution of the object point cloud, and explicitly calculate the motion axis in an online manner.
@@
    Closed-loop integration of interactive perception and online axis estimation. By continuously refining the motion axis estimation based on the updates of 3D point cloud, our approach improves the manipulation precision, and enables adaptive and precise control throughout the manipulation process, overcoming the limitations of open-loop methods.
@@ II-A Vision-based Robotic Manipulation
Visual perception is crucial for robotic manipulation, given its heavy reliance on visual modalities. Existing works adopt various inputs, with each modality exhibiting distinct advantages and limitations. While RGB images offer rich texture, they lack depth information crucial for grasping [6]. To address this, some methods combine RGB with depth maps [15, 16], though RGB-D data often suffers from noise on transparent or reflective surfaces [17].
Unlike approaches that switch across modalities [18, 19, 20], RGBManip [12] uses only multi-view RGB to estimate 6D object poses. However, its single end-effector camera can miss key interaction dynamics. Point-cloud-based methods, such as Where2Act [21] and SAGCI-System [22], excel at articulating action affordances, with RLAfford [16] emphasizing interaction dynamics. Flowbot3D [23] leverages motion flows to handle dynamic articulated objects.
Our approach strives to integrate the strengths of RGB-based methods and point-cloud-based techniques by tracking RGB inputs and segmenting the point clouds corresponding to the motion part of articulated objects. In order to address the limitation of open-loop methods, we integrate an additional RGB-D camera into the scene, specifically aimed at capturing the manipulation process.
@@
II-B 3D Point Cloud Segmentation Empowered by 2D Foundation Models
Segmentation techniques for 2D images have become highly sophisticated, with the Segment Anything Model (SAM) [24] as a prominent example. However, transferring these methods to 3D point clouds remains challenging, limiting their applicability in robotic manipulation.
DeepLabv3 [25] adapts 2D segmentation to 3D by projecting point clouds onto 2D planes [26], which is robust at scale but loses accuracy due to projection. To address this, SAM-based 3D approaches have emerged. For instance, Point-SAM [27] extends SAM to generate segmentation masks from point or mask prompts, while PartSLIP [28] and PartSLIP++ [29] leverage GLIP [30] and SAM for fine-grained part-level segmentation.
In dynamic scenes, video segmentation models such as SAM2 preserve motion cues more effectively. Our method builds on the RGBManip environment by integrating a single RGB-D camera and applying robust noise-handling. By masking the moving parts of the target objects in the point cloud, we capture motion information that better informs the manipulation policy.
@@ II-C Axis Estimation for Articulated Objects
However, previous studies on axis estimation are typically open-loop, which causes a loss of the interaction dynamics. To address this issue, [34] utilizing force/torque sensor measurements to estimate the motion direction and the orientation of the axis, inferring the type of joint without prior knowledge of the object’s kinematics. It employs closed-loop control relies on a velocity controller, which ensures that the end-effector moves at the desired velocity.
Our approach also employs closed-loop axis estimation. Unlike [34] which adopts the implicit estimation via kinematic models, we leverage the geometric properties of motion to explicitly compute the axis of the articulated object.
@@ III METHOD
We introduce a novel approach for manipulating articulated objects, guided by explicit axis estimation derived from SAM2-based tracking. As depicted in Fig. 2, our pipeline consists of 3 core modules: (1) Interactive Perception & Init-Manipulation Module, (2) Segmentation & Tracking Module and (3) Axis Estimation & Manipulation Module.
Initially, the robot applies any interactive perception methods to manipulate the articulated object, creating a slight displacement. Our method leverages the robust tracking capabilities of SAM2 to capture the 3D dynamics of articulated objects. After point cloud augmentation, we can obtain the Oriented Bounding Boxes (OBBs) and explicitly compute the precise axis information of the articulated object to drive the robotic arm for manipulation.
@@ III-A Interactive Perception & Init-Manipulation Module
The Interactive Perception & Init-Manipulation Module employs any interactive perception methods to grasp the handle and perform initiate manipulation to articulated objects, producing slight displacements for further analysis. Our method incorporates RGBManip as the default implementation for this module, chosen for its recent success in RGB-only interactive perception. It is worth noting that RGBManip employs the original SAM-based multi-view object pose estimation for detecting and grasping the handle.
As RGBManip initiates the manipulation of articulated objects, an additional camera periodically records dynamic RGB-D data. The captured RGB data will be leveraged for subsequent segmentation and tracking tasks, while the corresponding depth data facilitate the reconstruction of the 3D point cloud of the evolving dynamic scene.
III-B Segmentation & Tracking Module
We exploit Grounding DINO [13] on the initial RGB frame to generate an anchor box for the target object using text prompt, such as “cabinet”. The first RGB frame serves as the starting point for SAM2’s tracking process [14]. The anchor box, corresponding to the target object, is input into SAM2 as a box prompt, enabling it to continuously track the object, and provide a mask for it throughout each moment of the dynamic manipulation process. With the object’s mask, we can extract the corresponding portion of the 3D point cloud representing the object from the dynamic scene.
We filter the raw object point cloud $P=\{p_{i}\}_{i=1}^{n}$ obtained above to remove outliers, ensuring that the filtered point cloud represents the region of interest of articulated object. Specifically, a point $p\in P$ is retained if and only if
$|U^{\circ}(p,r)|\geq\epsilon,$ (1)
in which
$U^{\circ}(p,r)=\{q\mid q\in P,0<|p-q|\leq r\},$ (2)
and ||$\cdot$|| ⋅ | is the number of element in the set.
We then compute the oriented bounding box (OBB) for the entire articulated object on the initial frame. Subsequently, by subtracting this OBB from each frame’s point cloud and applying another round of noise reduction, we obtain the point cloud representing the moving parts of the articulated objects. For implementation details, the segmentation can be refined by removing the protruding handle from the cabinet’s OBB to obtain a tighter OBB of the cabinet’s body, leading to a more precise motion-part segmentation.
@@ III-C Axis Estimation & Manipulation Module
The Axis Estimation & Manipulation Module calculates the axis of motion based on the segmented point cloud representing the moving parts of the articulated object identified by the Segmentation & Tracking Module. Task-specific geometric priors play an essential role in achieving accurate axis estimation, i.e., for a prismatic joint, the moving parts translate along the axis, whereas for a revolute joint, the moving parts rotate around the axis. By leveraging these task-based priors, we are able to explicitly calculate the joint’s axis.
Specifically, consider a single action of robotic manipulation, where the OBBs of the initial and final point cloud of the motion components of the articulated object are denoted as $obb_{\text{st}}$ and $obb_{\text{ed}}$, with their centers designated as $O_{\text{st}}$ and $O_{\text{ed}}$ respectively.
  • Prismatic joint. The axis pivot is defined as $O_{\text{st}}$, and the axis direction is estimated as the direction $\overrightarrow{d}=\overrightarrow{O_{\text{st}}O_{\text{ed}}}$.
  • Revolute joint. The axis pivot point is derived by identifying the intersection of the mid-perpendiculars along the longer edges in the top-down view of $obb_{\text{st}}$ and $obb_{\text{ed}}$. With the axis point $P$ established, the axis direction is ascertained by evaluating the sign of the dot product $\overrightarrow{d}\cdot\overrightarrow{t}$, where $\overrightarrow{t}=\overrightarrow{O_{\text{st}}P}\times\overrightarrow{z}$, among which $\overrightarrow{z}$ represents the positive direction of the z-axis.
We propose a closed-loop axis estimation refinement method predicated on a straightforward observation: as the manipulation of the joint assembly increases, the point cloud of its moving component becomes progressively more amenable to accurate reconstruction. The enhancement in the fidelity of the point cloud data, in turn, furnishes the axis estimation process with inputs that are increasingly precise and reliable.
[…]